# Local-global compatibility and the exceptional zero conjecture for $\operatorname{GL}(3)$

### v1

@@ Abstract.
We prove the exceptional zero conjecture of Greenberg–Benois for $p$-ordinary regular algebraic cuspidal automorphic representations of $\operatorname{GL}_{3}(\mathbb{A})$ which are Steinberg at $p$. In particular, we obtain the first cases of this conjecture for non-essentially-self-dual RACARs of $\operatorname{GL}_{n}(\mathbb{A})$.
Our proof has two main parts. In Part 1, we use $p$-arithmetic cohomology to prove an exceptional zero formula using Gehrmann’s automorphic $\mathcal{L}$-invariant. In Part 2 we prove the equality of automorphic and Fontaine–Mazur $\mathcal{L}$-invariants. As one of the key ingredients for this, we establish local-global compatibility at $\ell=p$ for Galois representations attached to $p$-ordinary torsion classes for $\operatorname{GL}_{n}$, confirming a conjecture of Hansen in this setting. We prove this for all $n$ following the strategy in the “10-author paper”, and use the $n=3$ case to deduce the desired equality of $\mathcal{L}$-invariants.
@@ 1. Introduction
There are many conjectures connecting arithmetic information to analytic properties of $L$-functions, such as the Birch–Swinnerton-Dyer (BSD) and Bloch–Kato conjectures, which concern the relationship between the size of Mordell–Weil/Selmer groups and complex $L$-values. A profitable approach to studying this relationship has been to use $p$-adic methods to prove $p$-parts of these conjectures. For example, if $E/\mathbb{Q}$ is an elliptic curve with (ordinary) semistable reduction at $p$, then there is $p$-adic $L$-function $L_{p}(E,s)$ attached to $E$, related to the complex $L$-function via
@@
where $e_{p}(-)$ is Coates–Perrin-Riou’s modified Euler factor at $p$. Via the ($p$-adic) Iwasawa Main Conjecture [SU14, Ski16], the analytic properties of $L_{p}(E,s)$ can be related to arithmetic invariants for $E$. As a consequence, if $e_{p}(E,\mathbbm{1})\neq 0$, then (1.0.1) gives a non-trivial relation between the $p$-adic and complex values at $s=1$ which can be exploited to deduce the $p$-part of the BSD conjecture in analytic rank zero. If $E$ has split multiplicative reduction at $p$ however, then $e_{p}(E,\mathbbm{1})=0$; the $p$-adic $L$-function $L_{p}(E,s)$ has an exceptional zero at $s=1$. In this case, (1.0.1) is vacuous and the relationship between $L_{p}(E,s)$ and $L(E,1)$ is not clear; more work is therefore needed to obtain similar results towards the BSD conjecture.
In the 1980s [MTT86], Mazur, Tate, and Teitelbaum conjectured an ‘exceptional zero formula’:
@@
whenever $E$ has split multiplicative reduction at $p$. In particular, differentiating removes the exceptional zero, establishing the desired link between $L_{p}(E,s)$ and $L(E,1)$, whilst introducing the $\mathcal{L}$-invariant $\mathcal{L}_{E}$ of $E$ at $p$. This $\mathcal{L}$-invariant is known to have a variety of arithmetic descriptions. For example, $\mathcal{L}_{E}$ equals $\log_{p}(q)/\operatorname{ord}_{p}(q)$, where $q$ is the Tate period of $E/\mathbb{Q}_{p}$; it measures the position of the Hodge filtration in the associated $(\varphi,N)$-module at $p$; it has connections to $p$-adic local Langlands for $E$; and it describes infinitesimal data in a Hida family through $E$. In particular, (1.0.2) predicts a new, purely $p$-adic, connection between analysis and arithmetic. It also suggests the first derivative of $L_{p}(E,s)$ at $s=1$ and the complex $L$-value $L(E,1)$ should (conjecturally) carry the same arithmetic data.
@@
In this article, we consider the analogue of this conjecture for automorphic representations of $\operatorname{GL}_{3}(\mathbb{A})$ which are Steinberg at $p$ (see (1.1.1) below). In particular, we prove an exceptional zero formula for the $p$-adic $L$-function attached to such automorphic representations, as constructed in the work of Loeffler and the third author [LW21]. This provides the first example of the Greenberg–Benois conjecture for regular algebraic non-essentially-self-dual cuspidal automorphic representations of $\operatorname{GL}_{n}(\mathbb{A})$. One of the key ingredients is local-global compatibility at $\ell=p$ for the Galois representations associated with Hida families (and $p$-ordinary torsion classes) for $\operatorname{GL}_{3}$; we establish such a result for general $\operatorname{GL}_{n}$ following the strategy in [ACC+23, §5], which is of independent interest, and which we anticipate will have applications towards proving automorphy lifting theorems.
@@ 1.1. Set-up and statement of the main results
We now describe the main results of this article in more detail. Fix an isomorphism $\mathbb{C}\cong\overline{\mathbb{Q}}_{p}$ throughout. Let $n\geqslant 2$ and let $\pi$ be a regular algebraic cuspidal automorphic representation (RACAR) of $\operatorname{GL}_{n}(\mathbb{A})$. Without loss of generality, we assume that $\pi$ is unitary, and we let $\omega_{\pi}$ denote its central character (which is a Dirichlet character). We assume that $\omega_{\pi}(-1)=(-1)^{(n+1)/2}$ if $n$ is odd.
@@
associated with $\pi$, uniquely determined by the property that the trace of $\rho_{\pi}(\operatorname{Frob}_{\ell})$ matches with the eigenvalue of a certain Hecke operator acting on $\pi_{\ell}^{\operatorname{GL}_{n}(\mathbb{Z}_{\ell})}$ for all but finitely many $\ell\neq p$ (here, $\operatorname{Frob}_{\ell}$ denotes an arithmetic Frobenius at the prime $\ell$ and $L/\mathbb{Q}_{p}$ is a sufficiently large finite extension). We impose the following ramification assumption on $\rho_{\pi}$ at $p$:
Assumption.
Suppose that $\rho_{\pi}$ is “Steinberg at pp”, i.e., one has
@@
where $\omega$ denotes the $p$-adic cyclotomic character and $_{i}$ are non-crystalline extensions (for all $i=1,\dots,n-1$).
This assumption implies that $\pi$ has trivial cohomological weight and, by the work of [Hev23], that $\pi_{p}$ is isomorphic to the Steinberg representation of $\operatorname{GL}_{n}(\mathbb{Q}_{p})$. We note that the converse should also be true, namely if $\pi$ has trivial cohomological weight and $\pi_{p}$ is Steinberg at $p$, then the above assumption holds for $\rho_{\pi}$; however, this currently only known up to semisimplification (under some additional hypotheses).
We can quantify the extensions $_{i}$ by their associated $\mathcal{L}$-invariant. More precisely, by local class field theory we have an isomorphism $\operatorname{Hom}_{\operatorname{cts}}(\mathbb{Q}_{p}^{\times},L)\cong\operatorname{H}^{1}(\mathbb{Q}_{p},L)$, and we let $\mathcal{L}^{\operatorname{FM}}_{\pi,i}\in L$ denote the unique element such that
$\langle\operatorname{log}_{p}-\mathcal{L}^{\operatorname{FM}}_{\pi,i}\operatorname{ord}_{p},\;*_{(n-3)/2+i}\rangle=0$
where: $\langle-,-\rangle\colon\operatorname{H}^{1}(\mathbb{Q}_{p},L)\times\operatorname{H}^{1}(\mathbb{Q}_{p},L(1))\to L$ denotes the local Tate duality pairing and $\operatorname{log}_{p}$ (resp. $\operatorname{ord}_{p}$) denotes the $p$-adic logarithm (resp. $p$-adic valuation) normalised so that $\operatorname{log}_{p}(p)=0$ (resp. $\operatorname{ord}_{p}(p)=1$). Note that $\mathcal{L}^{\operatorname{FM}}_{\pi,i}$ exists by the assumption that $_{(n-3)/2+i}$ is non-crystalline.
@@
In particular $e_{p}^{(c)}(\pi_{p},\mathbbm{1},s)$ always has a simple zero at $s=c$, and hence $\mathscr{L}_{p}^{(c)}(\pi,\mathbbm{1})=0$; the $p$-adic $L$-function has an “exceptional zero” at the trivial character coming from $e_{p}(-)$. Assuming the existence of such a $p$-adic $L$-function, the Greenberg–Benois conjecture [Gre94, p.166] says the following:
@@
In particular, nothing is known about the exceptional zero conjecture for $\operatorname{GL}_{n}/\mathbb{Q}$ outside the essentially self-dual setting. The reason for this is all of the known exceptional zero results in the essentially self-dual setting exploit the fact that one can find $p$-adic families through $\pi$ with a Zariski dense set of classical specialisations – something which is not expected for non-essentially-self-dual representations (see [APS08, CM09] and Remark 1.1.3 below).
Beyond this, even constructing the Coates–Perrin-Riou $p$-adic $L$-function is difficult. Recently, in work of Loeffler–Williams [LW21], the $p$-adic $L$-function for $\pi$ was constructed for not necessarily essentially self-dual automorphic representations of $\operatorname{GL}_{3}(\mathbb{A})$. It is natural to ask whether one can prove an exceptional zero formula for this $p$-adic $L$-function. This is one of the main results of this article, which is the first case of an exceptional zero formula for non-essentially-self-dual regular algebraic cuspidal automorphic representations of $\operatorname{GL}_{n}(\mathbb{A})$:
@@ 1.2. Method of proof
We now specialise to the case $n=3$. The proof of Theorem A is naturally divided into two parts: an automorphic side and a Galois side. This is reflected in the structure of this article. More precisely, the first step (Theorem B) is to prove an exceptional zero formula replacing $\mathcal{L}_{\pi,c+1}^{\operatorname{FM}}$ with an “automorphic $\mathcal{L}$-invariant” $\mathcal{L}^{\operatorname{Aut}}_{\pi,c+1}$ (constructed in the work of Gehrmann [Geh21]). This part is completely unconditional and intrinsic to $\pi$ (and even independent of the existence of $\rho_{\pi}$). The second step is to then show that $\mathcal{L}^{\operatorname{Aut}}_{\pi,c+1}=\mathcal{L}^{\operatorname{FM}}_{\pi,c+1}$. The key input is to show that the family of Galois representations associated with the Hida family through $\pi$ satisfies local-global compatibility at $\ell=p$. In particular, we show this family is trianguline with the correct parameters (Theorem D), confirming cases of a conjecture of Hansen [Han17, Conj. 1.2.2].
@@ 1.2.1. The automorphic side
and let $\overline{B}$, $\overline{P}_{1}$, and $\overline{P}_{2}$ denote the opposites of these parabolic subgroups. Just as the Galois $\mathcal{L}$-invariant was defined in terms of extensions of powers of the $p$-adic cyclotomic character, the automorphic $\mathcal{L}$-invariant is defined in terms of certain extensions of (generalised) Steinberg representations.
@@ 1.3. Structure of the paper
In Part 2, we prove Theorems C and D. In §7 we first prove Theorem C in essentially self-dual cases via a comparison to Rosso’s work [Ros16, Ros15]. In §8, we study the $\operatorname{GL}(3)$-eigenvariety and Kozsul complexes of $p$-adic distributions, implementing the strategy of Gehrmann–Rosso [GR22] to prove Theorem C in general, up to a crucial trianguline hypothesis (Theorem 8.3.5). In §9, the most important section in Part 2, we prove Theorem D and show that it implies Theorem 8.3.5, thus completing the proof of Theorem C (and hence Theorem A).

### v2

@@ Abstract.
We prove exceptional zero conjectures for $p$-ordinary regular algebraic cuspidal automorphic representations of $\operatorname{GL}_{3}(\mathbb{A})$ which are Steinberg at $p$. We make no self-duality assumptions.
The paper has two parts. In Part 1, we use $p$-arithmetic cohomology to unconditionally prove an automorphic exceptional zero conjecture in this setting, using Gehrmann’s automorphic $\mathcal{L}$-invariant.
In Part 2 we prove, under mild assumptions that are expected to always hold, the equality of automorphic and Fontaine–Mazur $\mathcal{L}$-invariants, and thus deduce cases of the full Greenberg–Benois exceptional zero conjecture. As one of the key ingredients for this, we establish local-global compatibility at $\ell=p$ for Galois representations attached to $p$-ordinary torsion classes for $\operatorname{GL}_{n}$, confirming a conjecture of Hansen in this setting. We prove this for all $n$ following the strategy in the “10-author paper”, and use the $n=3$ case to deduce the desired equality of $\mathcal{L}$-invariants.
@@ 1. Introduction
There are many conjectures connecting arithmetic information to analytic properties of $L$-functions, such as the Birch–Swinnerton-Dyer (BSD) and Bloch–Kato conjectures, which concern the relationship between the size of Mordell–Weil/Selmer groups and complex $L$-values. A profitable approach to studying this relationship has been to use $p$-adic methods to prove $p$-parts of these conjectures. For example, if $E/\mathbb{Q}$ is an elliptic curve with (ordinary) semistable reduction at $p$, then there is a $p$-adic $L$-function $L_{p}(E,s)$ attached to $E$, related to the complex $L$-function via
@@
where $e_{p}(-)$ is Coates–Perrin-Riou’s modified Euler factor at $p$, and $L^{(p)}$ denotes the complex $L$-function with the Euler factor at $p$ removed. Via the ($p$-adic) Iwasawa Main Conjecture [SU14, Ski16], the analytic properties of $L_{p}(E,s)$ can be related to arithmetic invariants for $E$. As a consequence, if $e_{p}(E,\mathbbm{1})\neq 0$, then (1.0.1) gives a non-trivial relation between the $p$-adic and complex values at $s=1$ which can be exploited to deduce the $p$-part of the BSD conjecture in analytic rank zero. If $E$ has split multiplicative reduction at $p$, however, then $e_{p}(E,\mathbbm{1})=0$; the $p$-adic $L$-function $L_{p}(E,s)$ has an exceptional zero at $s=1$. In this case, (1.0.1) is vacuous and the relationship between $L_{p}(E,s)$ and $L(E,1)$ is not clear; more work is therefore needed to obtain similar results towards the BSD conjecture.
In the 1980s, Mazur, Tate, and Teitelbaum [MTT86] conjectured an ‘exceptional zero formula’
@@
whenever $E$ has split multiplicative reduction at $p$. In particular, differentiating removes the exceptional zero, establishing the desired link between $L_{p}(E,s)$ and $L(E,1)$, whilst introducing the $\mathcal{L}$-invariant $\mathcal{L}_{E}$ of $E$ at $p$. This $\mathcal{L}$-invariant is known to have a variety of arithmetic descriptions. For example, $\mathcal{L}_{E}$ equals $\log_{p}(q)/\operatorname{ord}_{p}(q)$, where $q$ is the Tate period of $E/\mathbb{Q}_{p}$; it measures the position of the Hodge filtration in the associated $(\varphi,N)$-module at $p$; it has connections to $p$-adic local Langlands for $E$; and it describes infinitesimal data in a Hida family through $E$. In particular, (1.0.2) predicts a new, purely $p$-adic, connection between analysis and arithmetic. It also suggests that the first derivative of $L_{p}(E,s)$ at $s=1$ and the complex $L$-value $L(E,1)$ should carry the same arithmetic data.
@@
In this article, we consider the analogue of this conjecture for automorphic representations of $\operatorname{GL}_{3}(\mathbb{A})$ which are Steinberg at $p$ (see Assumption 1.1.1 below). In particular, across Theorems A, B and C, we prove exceptional zero formulae for the $p$-adic $L$-functions attached to such automorphic representations, as constructed in the work of Loeffler and the third author [LW21]. This provides the first examples of the Greenberg–Benois exceptional zero conjecture for regular algebraic cuspidal automorphic representations of $\operatorname{GL}_{n}(\mathbb{A})$ without any self-duality assumption. One of the key ingredients is local-global compatibility at $\ell=p$ for the Galois representations associated with Hida families (and $p$-ordinary torsion classes) for $\operatorname{GL}_{3}$; we establish such a result for general $\operatorname{GL}_{n}$, following the strategy in [ACC+23, §5] (see Theorem D). This result is of independent interest, and analogous results have had applications towards proving automorphy lifting theorems.
@@ 1.1. Set-up and statement of the main results
We now describe the main results of this article in more detail. Fix an isomorphism $\mathbb{C}\cong\overline{\mathbb{Q}}_{p}$ throughout. Let $n\geqslant 2$ and let $\pi$ be a regular algebraic cuspidal automorphic representation (RACAR) of $\operatorname{GL}_{n}(\mathbb{A})$. Without loss of generality, we assume that $\pi$ is unitary, and we let $\omega_{\pi}$ denote its central character (which is a Dirichlet character). We assume that $\omega_{\pi}(-1)=(-1)^{(n+1)/2}$ if $n$ is odd (otherwise the complex $L$-function $L(\pi,s)$ has no critical values).
@@
associated with $\pi$, uniquely determined by the property that the trace of $\rho_{\pi}(\operatorname{Frob}_{\ell})$ matches with the eigenvalue of a certain Hecke operator acting on $\pi_{\ell}^{\operatorname{GL}_{n}(\mathbb{Z}_{\ell})}$ for all but finitely many $\ell\neq p$ (here, $\operatorname{Frob}_{\ell}$ denotes an arithmetic Frobenius at the prime $\ell$ and $L/\mathbb{Q}_{p}$ is a sufficiently large finite extension). We impose the following ramification assumptions on $\pi$ and $\rho_{\pi}$:
Assumption 1.1.1.
Suppose that:
  • (1)
    $\pi$ is $p$-ordinary and $\pi_{p}$ is isomorphic to the Steinberg representation of $\operatorname{GL}_{n}(\mathbb{Q}_{p})$;
  • (2)
    one has
@@
    where $\omega$ is the $p$-adic cyclotomic character and $_{i}$ are non-crystalline extensions (for all $i=1,\dots,n-1$).
Remark 1.1.3.
Under assumption (1), there conjecturally exist $p$-adic $L$-functions with exceptional zeros, without which the exceptional zero conjecture has no content. Note that (1) implies $\pi$ has trivial cohomological weight.
If (2) is satisfied and additionally $\overline{\rho}_{\pi}$ is absolutely irreducible and decomposed generic (see Remark 1.1.6 below), then by the work of Hevesi [Hev23], (1) is automatic. We note that the converse (1) $\Rightarrow$ (2) should also be true, however this is currently only known up to semisimplification (under the same irreducibility and decomposed generic assumptions; see Theorem 9.2.6).
We can quantify the extensions $_{i}$ by their associated Fontaine–Mazur $\mathcal{L}$-invariant. More precisely, by local class field theory we have an isomorphism $\operatorname{Hom}_{\operatorname{cts}}(\mathbb{Q}_{p}^{\times},L)\cong\operatorname{H}^{1}(\mathbb{Q}_{p},L)$, and we let $\mathcal{L}^{\operatorname{FM}}_{\pi,j}\in L$ denote the unique element such that
$\langle\operatorname{log}_{p}-\mathcal{L}^{\operatorname{FM}}_{\pi,j}\operatorname{ord}_{p},\;*_{(n-3)/2+j}\rangle=0$
where: $j\in\tfrac{1}{2}\mathbb{Z}$ is such that $(n-3)/2+j\in\mathbb{Z}\cap[1,n-1]$, $\langle-,-\rangle\colon\operatorname{H}^{1}(\mathbb{Q}_{p},L)\times\operatorname{H}^{1}(\mathbb{Q}_{p},L(1))\to L$ denotes the local Tate duality pairing and $\operatorname{log}_{p}$ (resp. $\operatorname{ord}_{p}$) denotes the $p$-adic logarithm (resp. $p$-adic valuation) normalised so that $\operatorname{log}_{p}(p)=0$ (resp. $\operatorname{ord}_{p}(p)=1$). Note that $\mathcal{L}^{\operatorname{FM}}_{\pi,j}$ exists by the assumption that $_{(n-3)/2+j}$ is non-crystalline.
@@
In particular $e_{p}^{(c)}(\pi_{p},\mathbbm{1},s)$ always has a simple zero at $s=c$, and hence $\mathscr{L}_{p}^{(c)}(\pi,\mathbbm{1})=0$; the $p$-adic $L$-function has an “exceptional zero” at the trivial character coming from $e_{p}(-)$. Assuming the existence of such a $p$-adic $L$-function, the Greenberg–Benois conjecture [Gre94, p.166] says:
@@
In particular, nothing is known about the exceptional zero conjecture for $\operatorname{GL}_{n}/\mathbb{Q}$ outside the essentially self-dual setting. One reason for this is that all of the known exceptional zero results described above exploit the fact that one can find $p$-adic families through $\pi$ with a Zariski dense set of classical specialisations – something which is not expected for non-essentially-self-dual representations (see [APS08, CM09] and Remark 1.1.6 below). Furthermore, all of the aforementioned results crucially use algebro-geometric/complex analytic methods, exploiting the fact that $\pi$ is a functorial transfer from a group whose locally symmetric spaces are Shimura varieties. Such methods cannot be applied for general $\pi$.
Beyond this, even constructing the Coates–Perrin-Riou $p$-adic $L$-function is difficult. Recently, in work of Loeffler–Williams [LW21], the $p$-adic $L$-function for $\pi$ was constructed for automorphic representations of $\operatorname{GL}_{3}(\mathbb{A})$ with no self-duality assumption. It is natural to ask whether one can prove an exceptional zero formula for this $p$-adic $L$-function. This is one of the main results of this article, which is the first case of an exceptional zero formula for non-essentially-self-dual regular algebraic cuspidal automorphic representations of $\operatorname{GL}_{n}(\mathbb{A})$:
@@
Our first major result towards this is Theorem B, where we prove an automorphic analogue of Conjecture 1.1.4. In particular, we prove the exact analogue of the stated exceptional zero formula, for both $L_{p}^{(0)}$ and $L_{p}^{(1)}$, with the Fontaine–Mazur $\mathcal{L}$-invariant $\mathcal{L}_{\pi,c+1}^{\operatorname{FM}}$ replaced by Gehrmann’s automorphic $\mathcal{L}$-invariant $\mathcal{L}_{\pi,c+1}^{\operatorname{Aut}}$ (from [Geh21]).
Theorem B is completely unconditional (beyond Assumption 1.1.1(1), which is needed for the statement to make sense), and is even independent of the existence of $\rho_{\pi}$. Assumption 1.1.1(2), and the other hypotheses in Theorem A, are needed only to prove the equality of $\mathcal{L}$-invariants (Theorem C).
Remark 1.1.6.
@@
Remark 1.1.8.
A prototype for this setting is that of Bianchi modular forms (RACARs for $\operatorname{GL}_{2}$ over an imaginary quadratic field). Here one encounters similar issues to those mentioned above: non-existence of $p$-adic families with a Zariski-dense set of classical specialisations, and underlying locally symmetric spaces that are not Shimura varieties. In this case, Conjecture 1.1.4 is not known. However, cases of the automorphic analogue of the conjecture were proved by the first and third authors in [BW19]. The methods used there (modular symbols) were very explicit; finding a more conceptual framework for obtaining such results, that could apply to higher rank settings, was a key aim of the present paper, and our approach is summarised in §1.2.1 below.
We hope to return to this case in future, using the methods of the present paper to prove cases of the full Greenberg–Benois exceptional zero conjecture for Bianchi modular forms.
@@ 1.2. Method of proof
We now specialise to the case $n=3$. The proof of Theorem A is naturally divided into two parts: an automorphic side (proving the automorphic exceptional zero formula, in Theorem B) and a Galois side (proving an equality of automorphic and Fontaine–Mazur $\mathcal{L}$-invariants, in Theorem C). This is reflected in the structure of this article.
The key input on the Galois side is to show that the family of Galois representations associated with the Hida family through $\pi$ satisfies local-global compatibility at $\ell=p$ (Theorem D). In particular, we show this family is trianguline with the correct parameters (Theorem D), confirming cases of a conjecture of Hansen [Han17, Conj. 1.2.2].
@@ 1.2.1. The automorphic side
and let $\overline{B}$, $\overline{P}_{1}$, and $\overline{P}_{2}$ denote the opposites of these parabolic subgroups. Just as the Galois $\mathcal{L}$-invariant was defined in terms of extensions of powers of the $p$-adic cyclotomic character, the automorphic $\mathcal{L}$-invariant is defined in terms of extensions of (generalised) Steinberg representations.
@@ 1.3. Structure of the paper
In Part 2, we prove Theorems C and D. In §7 we first prove Theorem C in essentially self-dual cases via a comparison to Rosso’s work [Ros16, Ros15]. In §8, we study the $\operatorname{GL}(3)$-eigenvariety and Koszul complexes of $p$-adic distributions, implementing the strategy of Gehrmann–Rosso [GR22] to prove Theorem C in general, up to a crucial trianguline hypothesis (Theorem 8.3.5). In §9, the most important section in Part 2, we prove Theorem D and show that it implies Theorem 8.3.5, thus completing the proof of Theorem C (and hence Theorem A).